Skip to content. | Skip to navigation

Personal tools
You are here: GES DISC Home Alerts OCO-2 Post-anomaly Data Status

Production and Distribution of OCO-2 science data resumes

warn_icon.gif IMPORTANT MESSAGE Sep 23, 2016    OCO-2 Post-anomaly Data Status

The OCO-2 instrument experienced an anomaly that interrupted routine operations at 4:28 UTC on August 10, 2016 (Orbit 11212). The cause of the anomaly was resolved, and the instrument recovery process was initiated on August 14 by conducting a standard, week-long decontamination cycle.  The decontamination cycle was completed, and the instrument’s optical bench and detectors were at their nominal operating temperatures on 21 August (Orbit 11376). An initial analysis of the data indicated that, while the instrument was working nominally for science observations, its on-board calibration mechanism had not initialized properly, and was out of alignment, precluding the routine production and delivery of calibrated science data products.  The calibrator alignment issue was diagnosed and resolved on 20:34 UTC on September 15, 2016.  After validating the quality of the data, the standard version 7 (v7) forward streams for L1B and L2 data were reinitiated on September 20, 2016 (Orbit 11815). 

There will be a permanent data gap from August 10 (Orbit 11212) to August 21 (Orbit 11376) because no science observations were collected during that period.  The L1B and L2 science data collected between August 21 and September 20, while the calibrator was not aligned properly, will not be processed as part of the forward v7 data stream, but will be processed as part of the August and September retrospective products (v7r).  These data will be processed using a non-standard calibration product. Users should report any issues that they encounter with these data.

Document Actions
NASA Logo -
NASA Privacy Policy and Important Notices
Last updated: Sep 29, 2016 01:14 PM ET